OpenStudy (anonymous):

PLEASE HELP!! Given the domain {-2,1,6} what is the range for the relation 4x+y=3? A.{-11, 1, 21} B.{11, -1, -21} C.{-1, 4, 9} D.{-5, 7, 27}

OpenStudy (anonymous):

@hartnn Do you think you can help me?

hartnn (hartnn):

just [lug in those values in 'x' and find the value of 'y' when x= -2, y=.... ?

hartnn (hartnn):

*plug

OpenStudy (anonymous):

how would i find y?

hartnn (hartnn):

4x+y=3 4 (-2)+y = 3 -8+y=3 y= 3+8 =11 did you get this ? can you now find 'y' for x=1 and x=6 ?

OpenStudy (anonymous):

I sort of get it but I don't get it. Where did the 8 come from?

hartnn (hartnn):

4*2 =8 so -4*2 =-8

OpenStudy (anonymous):

is it -5

OpenStudy (anonymous):

3+-8=-5

hartnn (hartnn):

no, it'll be 3+8 because we ADD 8 on both sides to get rid of -8 on left. 3+8=11

OpenStudy (anonymous):

Ohh okay I see

OpenStudy (anonymous):

What comes next

hartnn (hartnn):

can you now find 'y' for x=1 and x=6 ?

hartnn (hartnn):

also, if you look at choices, the correct one is identifiable now.

OpenStudy (anonymous):

11

OpenStudy (anonymous):

4(-2)+11=3

hartnn (hartnn):

we already found 1 value of y as 11 so what was the point of doing that ?

OpenStudy (anonymous):

Wait so the answer is B

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Okay. Thanks

hartnn (hartnn):

welcome.